OpenStudy (anonymous):

Simplify the expression \[\frac{7}{x+2}-\frac{3x}{(x+2)^2}-\frac{5}{x}\]

OpenStudy (anonymous):

first take lcm of denominators

OpenStudy (anonymous):

\[\frac{7x(x + 2) - 3x.x - 5(x+2)^2}{x(x+2)^2}\] Solve it now??

OpenStudy (anonymous):

Use this also in solving: \[(a + b)^2 = a^2 + b^2 + 2ab\]

OpenStudy (anonymous):

I got the same thing so far @waterineyes

OpenStudy (anonymous):

But the book has \[\frac{5x^2-6x-20}{x(x+2)^2}\] and I got \[\frac{-x^2-6x-20}{x(x+2)^2}\]

OpenStudy (anonymous):

Can you scan and attach the book page where this question is??

OpenStudy (anonymous):

The book has a mistake there will come + instead of - in front of second term..

OpenStudy (anonymous):

I see what I did... it's suppose to be: \[\frac{7}{x+2}+\frac{3x}{(x+2)^2}-\frac{5}{x}\]

OpenStudy (anonymous):

Yes this will lead to your answer...

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

Haha, yeah. Thanks @waterineyes!!! and @problemsolver!!!